OpenStudy (anonymous):

Solve the following proportion: 4/a+4=a/8

OpenStudy (anonymous):

I cross multiply first, right? @ganeshie8

ganeshie8 (ganeshie8):

\(\large \dfrac{4}{a} + 4 = \dfrac{a}{8}\)

ganeshie8 (ganeshie8):

like this ?

OpenStudy (anonymous):

no, a+4 is under the 4 \[\frac{ 4 }{ a+4 }=\frac{ a }{ 8 }\]

ganeshie8 (ganeshie8):

Good :) then you're right ! yes you cross multiply first

OpenStudy (anonymous):

\[\frac{ a(a+4) }{ 32 }\] \[\frac{ a^2+4a}{ 32 }\]

OpenStudy (texaschic101):

a(a + 4) = (8)(4) a^2 + 4a = 32 continue...

OpenStudy (anonymous):

do I factor out the numerator, i think that's what it's called. (top)

OpenStudy (texaschic101):

a^2 + 4a - 32 = 0 I believe your gonna have to use quadratic formula

OpenStudy (texaschic101):

wait....I think you can factor that (a + 8)(a - 4) = 0

OpenStudy (anonymous):

ooooh ok, now I remember! Yea, like that. The answers are: a=-8 & a=4

OpenStudy (texaschic101):

yep...thats what I got

OpenStudy (anonymous):

Yaaaaay!! I solved the next one, I just want to make sure it's correct.